.Speaking of bacon, demand for bacon can be represented as P = 65 – 0.02Q. The constant private marginal cost for the firms in the market producing the bacon is $12 per unit. As it turns out, the industry creates runoff from the pig operations causing $7 in external damages to the area for every unit of bacon produced

c. Suppose that this market was not highly competitive, but in fact, was a monopoly market. If the firm was maximizing its profit, how much would it product? What would the price be for bacon? (Assume the monopolist won’t compensate for any external damage caused by its production.) Q = ____________________ P = _____________________

d. How does the monopoly outcome compare (part c) to the socially optimal outcome (part b)?

e. Is one outcome (from part d) clearly preferred over the other? Why or why not?
Answer #1

The Monopoly can be defined as a market struture when there are plenty of buyers but there is only one seller for the product.

The demand for the bacon is given below:

P=65-0.02Q

The total revenue will be P*Q=(65-0.02Q)*Q=65Q-0.02Q2

The Marginal revenue will be MR=65-0.04Q

c. If the Mopolist doesn't pay for the external damage cost by the production, the marginal cost for the monopolist is MC=$12. To maximize the profit, he will comapre the MR to MC and get the equilibrium quantity and Price as computed below:

MR=MC

65-0.04Q=$12

53=0.04Q

Q*=1,325

P*=65-0.02Q=38.5

The profit earned by the monopolist is TR-TC=(38.5-12)*1325=35,112.5

d. For socially optimal output , the marginal damage caused by thr production should be taken inot account, the MR should be equated to Social marginal cost(SMC)

MR=SMC

65-0.04Q=$19

46=0.04Q

Q*=1,150

P*=65-0.02Q=42

The profit earned by the monopolist is TR-TC=(42-19)*1325=26,450

e. As prices increases and quantity decreases in part (d) then in part(c), the consumer and producer surplus both have fallen . Therefore, outcome from part (d) is not prefered over outcome from part (c).